A particular first-order reaction has a rate constant of 1.55x10^2 s^-1 at 35.0 degrees Celsius. What is the magnitude of k at 95.0 degrees Celsius if Ea=55.5 kJ/mol?


1
Expert's answer
2021-03-10T17:51:50-0500

lnK1/K2= Ea/R (1/T1 - 1/T2)

ln155/K2= 55.5x103/8.314 (1/35-1/95)

ln155/K2= 6675.489(0.018)

ln155/K2= 120.46

155/K2= e120.46

K2= 155/2.0x1052

K2= 7.50 x 10-51